d Use logarithmic differentiation to find dz . (1703-944) (22 – 3)(23 + 4) 25(2 – 1)
Answer #1

(22-3)(22+4) f(2)= 25(2-1) take logrithen of both sides, In{f(z)3 = en [22-3) (2²+4) 25 (2-1) = In (z23) + ln (22+4) – en z e